OpenStudy (anonymous):

Write the general form of the equation which matches the graph below. In complete sentences, explain the process taken to find this equation.

OpenStudy (anonymous):

thats the graph

OpenStudy (amistre64):

parabola eh; y^2 = 4ax if i recall correctly

OpenStudy (amistre64):

that will of course have to be modified to this particular graph

OpenStudy (amistre64):

a is equal to the value of x at that dotted line; cant tell what it is from here tho

OpenStudy (amistre64):

and since x is always negative; that means its gonna be a -y^2

OpenStudy (anonymous):

it says i have to put it in x=a(y-h)^2-k format

OpenStudy (amistre64):

right, but lets get this format right before we try to shift things about

OpenStudy (amistre64):

-y^2 = 4ax x = -\(\cfrac{1}{4a}\) (y-3)^2 is what it looks like to me

OpenStudy (amistre64):

i cant tell the value for a; what does the dotted line cross at?

OpenStudy (anonymous):

it crosses at 4,0

OpenStudy (amistre64):

if that right, then that denom is 16

OpenStudy (anonymous):

so the final equation is x=1/16(y-3)^2?

OpenStudy (amistre64):

id say yes; except for make it a negative on your fraction

OpenStudy (amistre64):

x is always negative in this graph so you have to put it in there

OpenStudy (anonymous):

okay thank you so much

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

thnx

OpenStudy (amistre64):

2 years ago eh ... and sure, youre welcome :)

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

you wanted to know how we get: x=-1/4a(y-3)^2

OpenStudy (amistre64):

the way i find these has changed over the years

OpenStudy (amistre64):

since the graph opens to the left, then we know this is going to have to be part of some x = -y^2 setup but when y=3, we want x to be 0; so we modify it as: x = (y-3)^2 the only thing left to do is to determine how to spread it out, in general, the point directly across from the focus will help us out